OpenStudy (anonymous):

Find a so that f(x)= {ax^2-9 if x 2 is continuous at x=2

OpenStudy (anonymous):

this is much easier than it seems. replace \(x\) by 2 in both expressons, set them equal, solve for \(a\)

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} ax^2-9& \text{if} & x \leq 2 \\ 4x+3& \text{if} & x > 2 \end{array} \right. \]

OpenStudy (anonymous):

you get \[a\times 2^2-9=4\times 2+3\] \[4a-9=11\] etc

OpenStudy (anonymous):

but don't i need to set both to a limit ?

OpenStudy (anonymous):

yes of course, but both of these are continuous functions, so the limit is what you get when you replace \(x\) by 2

OpenStudy (anonymous):

the idea is that you are trying to match them up so they are equal where the function changes definition

OpenStudy (anonymous):

in other words, they should agree when \(x=2\) when \(x=2\) the first one is \(4a-9\) and you want to make sure it agrees with the second one, which would be \(11\)

OpenStudy (anonymous):

i got a = 9/4

OpenStudy (anonymous):

you will notice that the first thing is said was "it is easier than it seems" it really is just elementary algebra right?

OpenStudy (anonymous):

i don't think that is right did you solve \[4a-9=11\] ?

OpenStudy (anonymous):

i don't understand why you are making 4a-9 equal to 11

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

when you replace \(x\) by 2 in the top formula \(ax^2-9\) what do you get?

OpenStudy (anonymous):

4a-9

OpenStudy (anonymous):

ok good. now you know that \(ax^2-9\) is a polynomial, so the limit is identical to the value of the function at any number because all polynomials are continuous everwhere

OpenStudy (anonymous):

so the value at \(x=2\) is \(4a-9\) now what do you get when you replace \(x\) by 2 in the second formula?

OpenStudy (anonymous):

so the lim as x->2- ax^2-9 = 2?

OpenStudy (anonymous):

in the seconand formula it = 11

OpenStudy (anonymous):

no not 2, but rather \(f(2)\)

OpenStudy (anonymous):

the limit is the same as the value of the function, not the same as \(x\)

OpenStudy (anonymous):

\[\lim_{x\to 2}ax^2-9=a\times 2^2-9=4a-9\]

OpenStudy (anonymous):

so \[\lim x \rightarrow0 of ax ^{2}-9=f(2)\]

OpenStudy (anonymous):

and \(\lim_{x\to 2}4x+3=4\times 2+3=11\)

OpenStudy (anonymous):

right, thats what i have.

OpenStudy (anonymous):

ok so the limit from the left is \(4a-9\) and the limit from the right is \(11\) and you want the function to be continuous, so you have to make sure those limits are the same. ergo set \[4a-9=11\] and solve for \(a\)

OpenStudy (anonymous):

a=5

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so f(x)=11=f(2)=a ?

OpenStudy (anonymous):

now lets go back to the beginning as see how easy this was and what needed to be done replace \(x\) by 2 in both expressions, set them equal, solve for \(a\)

OpenStudy (anonymous):

what is \(a\) ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

so what is \(ax^2-9\) ?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

when u sub in 5

OpenStudy (anonymous):

no, but rather when you replace \(a\) by 5 in the expression \(ax^2-9\) ?

OpenStudy (anonymous):

i think there is some confusion here. \(a\) is just some constant, the coefficient of \(x^2\) for the first formula now you know \(a=5\) because if \(a=5\) then the top formula becomes \(5x^2-9\) and the bottom formula is \(4x+3\) and if you replace \(x\) by 2 in either formula, you get \(11\) in both cases, meaning the limit is the same from the left and from the right you answer is \[f(x) = \left\{\begin{array}{rcc} 5x^2-9& \text{if} & x \leq 2 \\ 4x+3& \text{if} & x > 2 \end{array} \right.\]

OpenStudy (anonymous):

you can see from the answer i wrote, that if \(x=2\) the first one gives \(5\times 4-9=11\) and the second one gives \(4\times 2+3=11\) so both limits are the same, and the function is continuous now

OpenStudy (anonymous):

so i have to re-write my function with the 5 instead of the a?

OpenStudy (anonymous):

yes, the whole question was to find out what \(a\) should be

OpenStudy (anonymous):

if \(a\) was not 5, say \(a=3\) then the function would not be continuous at 2, because the limit would not be 11, but it would still be 11 for the second formula

OpenStudy (anonymous):

okay thank-you,

OpenStudy (anonymous):

yw